The viscosity of a liquid detergent is supposed to average 800 centistokes at 25° C. A random sample of 16 batches of detergent is collected, and the average viscosity is 812. Suppose we know that the standard deviation of viscosity is σ = 25 centistokes: (a) State the hypotheses that should be tested, (b) Test these hypotheses using a = 0.05. What are your conclusions?, (c) What is the P-value for the test?, (d) Find a 95 percent confidence interval on the mean.   
Answer #1

To, test whether the average viscocity is 800 centistokes at the 5% level of significance or not using the z-test at the leve

(c) The calculation of the P-value: Step6: The P-value approach: 2P(2>2 ) = 2(1-P(z<zo)) =2(1-P(Z <1.92)) =2(1-0.972571) =2(0

The sample size is n=16. The sample mean is X = 812. The sample standard deviation is o = 25. The formula for calculating the